OpenStudy (anonymous):

find a formula for the inverse f(x)=x+6

OpenStudy (anonymous):

so y=x+6, replace x with y, and y with x so you have x=y+ 6 now solve for y

OpenStudy (anonymous):

subtract 6 from both sides and you have y= x -6